A regular octagon of side 10 is inscribed in a circle. Find the radius of the circle.



1
Expert's answer
2021-11-01T07:07:48-0400


"\alpha=360\degree/8=45\degree"


The Law of Cosines



"10^2=r^2+r^2-2r^2\cos45\degree""r^2=\dfrac{100}{2-\sqrt{2}}""r=10\sqrt{1+\dfrac{\sqrt{2}}{2}}"